Calculate the wavelength in metres associated with an 80g mass moving at 3km/hr
1
Expert's answer
2019-04-01T07:48:06-0400

The de Broglie wavelength is the wavelength, λ, associated with a massive particle and is related to its momentum, p, through the Planck constant, h (6.62607015×10−34 J⋅s or kg⋅m2/s):

{\displaystyle \lambda ={\frac {h}{p}}={\frac {h}{mv}}.}λ,= h/mv


 

80 g = 0,08 kg

3 km/h = 3000 m/h = 0,8333 m/s

 


λ,= 6.626×10−34 /(0,08 kg*0,8333 m/s) =9,943×10−33 m